OpenStudy (anonymous):

1/x-1divided by 1/x^2-1 the lim x->0

OpenStudy (shamim):

is it\[\lim_{x \rightarrow 0}\frac{ \frac{ 1 }{ x-1 } }{ \frac{ 1 }{ x ^{2}-1 } }\]

OpenStudy (anonymous):

yes

OpenStudy (shamim):

ok

OpenStudy (shamim):

\[\lim_{x \rightarrow 0}\frac{ 1 }{ x-1 } \times (x-1)(x+1)=\lim_{x \rightarrow 0}x+1=1\]

OpenStudy (shamim):

medal please

OpenStudy (anonymous):

thanks

OpenStudy (shamim):

welcome